Sol Summary – November 16th

MDRS Daily Summary Report for Sol 13.

Summary Title: Organizing MDRS.

Mission Status: Preparing to leave MDRS. We did some improvements in MDRS fixing some issues in the physical facilities.

Sol Activity Summary: We just concluded the cleaning procedures of the entire installation, also cleaning the GreenHab and the laboratory. We developed many different activities organizing and fixing the facilities of MDRS.

Look Ahead Plan: Tomorrow, probably will happen an EVA, but We are just waiting the SurvivorMan to decide about would be interesting to him.

Anomalies in work: Nope.

Weather: 17o C, 62o F.

Crew Physical Status: Fine.

EVA: We planed an EVA today, but not was possible. We planed to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives.


Reports to be file:

Commanders Report

Daily Summary Report

Operation Report

GreenHab Report

HSO Report

XO Report

Crew-photos Crew 182 16Nov2017.


Support Requested: Nope.

Copyright © The Mars Society. All rights reserved. | Main Site